# Channel Measurement, Modeling, and Simulation for 6G: A Survey and Tutorial

### v3

@@ II-A Large-Scale Channel Fading
Large-scale fading comes from path loss as a function of distance and shadowing by large objects such as buildings and hills [58]. The path loss represents the reduction in radio wave power over a long distance [59]. Shadow fading is mainly due to the multipath blockage, which also causes significant attenuation of the signal strength. The 5G standard provides three types of path loss models, including the close-in (CI) model, floating-intercept (FI) model, and alpha-beta-gamma (ABG) model [61, 62, 63]. The FI model is expressed as
@@
where $\textrm{PL}_{\textrm{FI}}(d)$ denotes the path loss in dB, which is a function of the Tx-Rx separation distance $d$. $\beta_{\textrm{FI}}$ and $\alpha_{\textrm{FI}}$ are two fitted parameters. $X_{\sigma}^{\textrm{FI}}$ is a zero-mean Gaussian variable with standard deviation $\sigma$ in dB representing shadow fading. Note that $\beta_{\textrm{FI}}$ is frequency dependent.
The CI path loss model can be provided by (2).
@@
where $f$ is the center frequency, $d$ is the Tx-Rx separation distance greater than the reference distance $d_{0}$, c denotes the speed of light, $\alpha_{\textrm{CI}}$ is the single model parameter known as the path loss exponent (PLE), and $X_{\sigma}^{\textrm{CI}}$ is the zero-mean Gaussian variable, which describes the large-scale signal fluctuations about the mean path loss over distance.
The third type of path loss model, ABG model is represented as a function of frequency and distance [62, 64]. It can describe the path loss at various frequencies, and is expressed as
@@
where $d_{0}$ is the received power reference point, typically set to 1 m, here $f$ is in GHz. $\alpha_{\textrm{ABG}}$, $\beta_{\textrm{ABG}}$ and $\gamma_{\textrm{ABG}}$ are fitting parameters, where $\alpha_{\textrm{ABG}}$ and $\gamma_{\textrm{ABG}}$ characterize the distance dependence and frequency dependence of path loss, respectively, and $\beta_{\textrm{ABG}}$ is an offset value in dB. The typical value of $\gamma_{\textrm{ABG}}$ is 2, which corresponds to free space propagation. $X_{\sigma}^{\textrm{ABG}}$ is the zero-mean Gaussian variable. In the ABG model, the path loss varies with the frequency and distance. In [65], it was demonstrated that the CI and ABG models offer comparable modeling performance upon using field measurement data.
@@ II-B Small-Scale Channel Fading
Geometry-based statistical channel modeling is the mainstream approach adopted by both ITU and 3GPP in 4G and 5G. According to the definition of the geometry-based stochastic model (GBSM) [66], the channels are generated by summing the contributions of rays based on geometry principles and specific small-scale statistical channel parameters from measurement. Geometry-based modeling enables the channel to be independent of the antenna and allows the configuration of different antennas and different element patterns. Additionally, for individual channel snapshots, the channel parameters, such as delay, power, angle of arrival, and angle of departure, are stochastically determined based on statistical distributions extracted from channel measurements. In the GBSM-based channel model of the 5G standard, a new dimension of elevation are introduced at both Tx and Rx, enabling support for 3D MIMO.
The 3D GBSM based channel model is illustrated in Fig. 3. The small sphere with several dots inside represents a scattering region that generates a cluster of propagation paths. Each cluster is constituted by $M$ rays, and $N$ clusters are assumed. Tx and Rx are equipped with $S$ and $U$ antennas, respectively. The small-scale parameters (SSPs), such as the delay $\tau_{n,m}$, zenith angle of arrival $\theta_{n,m,\textrm{ZOA}}$, azimuth angle of arrival (AOA) $\phi_{n,m,\textrm{AOA}}$, azimuth angle of departure (AOD) $\phi_{n,m,\textrm{AOD}}$, and zenith angle of departure (ZOD) $\theta_{n,m,\textrm{ZOD}}$, are assumed to be different for the $m$th ray in the $n$th cluster. Radio waves propagate in 3D space, and MPCs exhibit not only horizontal-dimension characteristics but also vertical-dimension characteristics.
Figure 3: Illustration of the 3D GBSM-based channel model [67].
@@
where $a_{n,m}$ is the complex amplitude of the $m$th MPC within the $n$th cluster, $F_{\textrm{Rx},u}$ and $F_{\textrm{Tx},s}$ are the $u$th antenna pattern at the Rx side and the $s$th antenna pattern at the Tx side [68, 69], respectively. $\hat{r}_{\textrm{Rx},n,m}$ is the unit vector in the spherical coordinate system at the Rx side, $\vec{d}_{\textrm{Rx},u}$ is the location vector of the $u$th antenna at the Rx side, and the subscript Tx similarly represents the Tx side. $\delta(\cdot)$ is the Dirac function, $\lambda$ is the wavelength, and $M_{n}$ is the number of delay bins. $\bm{\Omega_{n,m}}$ and $\bm{\Psi_{n,m}}$ are the angles of arrival and departure of the MPC, respectively. They include both the elevation and azimuth angles, $\theta$ and $\phi$, for 3D MIMO systems as follows:
@@
In practical 5G systems, antennas can be designed and configured to exhibit specific radiation patterns that enhance the system performance, and slant polarized antennas are usually used. Thus, the radiation patterns need be modeled when generating the CIR. Horizontal and vertical linear polarizations and the $+45^{\circ}/-45^{\circ}$ slant polarization are examples of orthogonal polarizations [70]. When dual polarization is utilized, (4) is changed to (6).

### v4

@@ II-A Large-Scale Channel Fading
Large-scale fading comes from path loss as a function of distance and shadowing by large objects such as buildings and hills [58]. Path loss represents the reduction in radio wave power over a long distance [59]. Shadow fading is due mainly to multipath blockage, which can also cause a significant attenuation of signal strength. The 5G standard provides three types of path loss models, including the close-in (CI) model, floating-intercept (FI) model, and alpha–beta–gamma (ABG) model [61, 62, 63]. The FI model is expressed as
@@
where $\textrm{PL}_{\textrm{FI}}(d)$ denotes the path loss in dB, which is a function of the Tx–Rx separation distance $d$. $\beta_{\textrm{FI}}$ and $\alpha_{\textrm{FI}}$ are two fitted parameters. $X_{\sigma}^{\textrm{FI}}$ is a zero-mean Gaussian variable with a standard deviation of $\sigma$ in dB, representing shadow fading. Note that $\beta_{\textrm{FI}}$ is frequency dependent.
The CI path loss model can be provided by (2)
@@
where $f$ is the center frequency, $d$ is the Tx–Rx separation distance greater than the reference distance $d_{0}$, c denotes the speed of light, $\alpha_{\textrm{CI}}$ is the single model parameter known as the path loss exponent (PLE), and $X_{\sigma}^{\textrm{CI}}$ is the zero-mean Gaussian variable, which describes the large-scale signal fluctuations about the mean path loss over distance.
The third type of path loss model, the ABG model, is represented as a function of frequency and distance [62, 64]. It can describe the path loss at various frequencies and is expressed as
@@
where $d_{0}$ is the received power reference point, typically set to 1 m; here, $f$ is in GHz. $\alpha_{\textrm{ABG}}$, $\beta_{\textrm{ABG}}$ and $\gamma_{\textrm{ABG}}$ are fitting parameters, where $\alpha_{\textrm{ABG}}$ and $\gamma_{\textrm{ABG}}$ characterize the distance dependence and frequency dependence of path loss, respectively, and where $\beta_{\textrm{ABG}}$ is an offset value in dB. The typical value of $\gamma_{\textrm{ABG}}$ is 2, which corresponds to free space propagation. $X_{\sigma}^{\textrm{ABG}}$ is the zero-mean Gaussian variable. In the ABG model, path loss varies with frequency and distance. [65] demonstrated that the CI and ABG models offer comparable modeling performance when field measurement data are used.
@@ II-B Small-Scale Channel Fading
Geometry-based statistical channel modeling is the mainstream approach adopted by both the ITU and 3GPP in 4G and 5G. According to the definition of the geometry-based stochastic model (GBSM) [66], the channels are generated by summing the contributions of rays on the basis of geometry principles and specific small-scale statistical channel parameters from measurements. Geometry-based modeling enables the channel to be independent of the antenna and allows the configuration of different antennas and different element patterns. Additionally, for individual channel snapshots, the channel parameters, such as delay, power, angle of arrival, and angle of departure, are stochastically determined on the basis of statistical distributions extracted from channel measurements. In the GBSM-based channel model of the 5G standard, a new dimension of elevation is introduced at both Tx and Rx, enabling support for 3D MIMO configurations.
The 3D GBSM-based channel model is illustrated in Fig. 3. The small sphere with several dots inside represents a scattering region that generates a cluster of propagation paths. Each cluster consists of $M$ rays, and $N$ clusters are assumed. The Tx and Rx are equipped with SS and $U$ antennas, respectively. The small-scale parameters (SSPs), such as the delay $\tau_{n,m}$, zenith angle of arrival $\theta_{n,m,\textrm{ZOA}}$, azimuth angle of arrival (AOA) $\phi_{n,m,\textrm{AOA}}$, azimuth angle of departure (AOD) $\phi_{n,m,\textrm{AOD}}$, and zenith angle of departure (ZOD) $\theta_{n,m,\textrm{ZOD}}$, are assumed to differ for the $m$th ray in the $n$th cluster. Radio waves propagate in 3D space, and MPCs exhibit not only horizontal-dimension characteristics but also vertical-dimension characteristics.
@@
where $a_{n,m}$ is the complex amplitude of the $m$th MPC within the $n$th cluster and where $F_{\textrm{Rx},u}$ and $F_{\textrm{Tx},s}$ are the $u$th antenna patterns at the Rx side and the $s$th antenna pattern at the Tx side [67, 68], respectively. $\hat{r}_{\textrm{Rx},n,m}$ is the unit vector in the spherical coordinate system at the Rx side, $\vec{d}_{\textrm{Rx},u}$ is the location vector of the $u$th antenna at the Rx side, and the subscript Tx similarly represents the Tx side. $\delta(\cdot)$ is the Dirac function, $\lambda$ is the wavelength, and $M_{n}$ is the number of delay bins. $\bm{\Omega_{n,m}}$ and $\bm{\Psi_{n,m}}$ are the angles of arrival and departure of the MPC, respectively. They include both the elevation and azimuth angles, $\theta$ and $\phi$, for 3D MIMO systems as follows:
@@
Figure 3: Illustration of the 3D GBSM-based channel model [69].
In practical 5G systems, antennas can be designed and configured to exhibit specific radiation patterns that enhance system performance, and slant polarized antennas are typically used. Thus, the radiation patterns need to be modeled when generating the CIR. Horizontal and vertical linear polarizations and the $45^{\circ}/-45^{\circ}$ slant polarization are examples of orthogonal polarizations [70]. When dual polarization is utilized, (4) is changed to (6).